OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 3 points) [2.02] MA.8.A.6.4 Which Property of Equality should be used first to solve the 8 + p/2=16 equation for p? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

maybe do some work yourself?

OpenStudy (anonymous):

@JamesWolf Maybe... Stay Off My Stuff?

OpenStudy (anonymous):

Subtraction Property of Equality 8 + p/2=16 Because first we will remove 8 by subtracting 8 from both the sides 8-8+p/ 2=16-8 p/2=8 Then we will use Multiplication Property of Equality beacuse we will multiply 2 with 8 p/2=8 p=8x2 p=16

OpenStudy (anonymous):

I hope I'm clear :) and @Zackeria_x3 leave don't fight

OpenStudy (anonymous):

Ok.. Thanks again ! :)

OpenStudy (anonymous):

@xoya :)

OpenStudy (anonymous):

So, the answer to this question is Subtraction Property of Equality :)

OpenStudy (anonymous):

Yes say? :)

OpenStudy (anonymous):

Yes, Thanks A-LOT :)

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

Your welcome :) Yeah yeah say?